Assume a muscle has a PCSA = 20 cm2 and Lo = 12 cm. Assume it is composed of 40% slow twitch fibers and 60% fast twitch fibers. Find the coefficients a and b for Hill’s equation.
Answer #1

The Hill equation is described as:
V = b (F0-F) / (F + a),
Where Fo is the maximum PCSA force given in the exercise.
To predict the different speeds of muscle contraction by varying the load, where V = contraction speed, F0 is the maximum force that the muscle is capable of expressing, in an isometric contraction, performed with an optimum degree of pre-stretching, F is instantaneous force in a certain place of the graph, a is constant with units of force, b is constant with units of speed. To determine a and b, curves must be constructed and data is missing. The most that is known is that in the slow fibers a = 0.25 * 1250 N and b = 0.25 * 72 cm / s, while in the fast a = 0.25 * 1250 N and b = 0.25 * 192 cm / s. in the case of intermediate percentages as they indicate it can not be predicted since there is a lack of data.